A group of 40 college students from a certain liberal arts college were randomly sampled and asked about the number of alcoholic drinks they have in a typical week. The purpose of this study was to compare the drinking habits of the students at the college to the drinking habits of college students in general. In particular, the dean of students, who initiated this study, would like to check whether the mean number of alcoholic drinks that students at his college in a typical week differs from the mean of U.S. college students in general, which is estimated to be 4.73. The group of 40 students in the study reported an average of 4.23 drinks per with a standard deviation of 3.85 drinks. Find the p-value for the hypothesis test. The p-value should be rounded to 4-decimal places. On excel please
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

n = 40

Mean pop = 4.73

sample mean = 4.23

Stdev = 3.85

Lets calculate the p-value.

read this statement "The purpose of this study was to compare the drinking habits of the students at the college to the drinking habits of college students in general. In particular, the dean of students, who initiated this study, would like to check whether the mean number of alcoholic drinks that students at his college in a typical week differs from the mean of U.S. college students in general," This means that we have to perform a 2 tailed test.

Therefore,

Ho: p1 = p2

Ha: p1 != p2

Lets calculate the Test statistic ( We use t-test since population deviation is unknown)

t = (xbar - Mu)/(s/sqrt(n)) = (4.23 - 4.73)/(3.85/sqrt(40)) = - 0.8214

So, the p-value is basically = 2*P(|t| > 0.8214 ) = 0.4164

In excel we have used this formula to calculate the p-value = 2*T.DIST(-0.8214, 39, TRUE)

We have multiplied by 2 because we have 2 addup p-values of both ends of the tails.